• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name V. Sulkosky

    Log entry time 21:22:41 on May16,2007

    Entry number 202348

    keyword=BCM Monitor and Current to 80 uA

    Over the past two hours the beam has been very trippy (about 23 trips
    in the past hour). The MCC operator said that the separation into
    Hall A had changed causing the problem. The trips started around the
    same time that Hall C started running 80 uA. After conferring with
    the spokespersons, we have decided to drop to the current to 80 uA to
    check if the trip rate improves. So far so good.

    FIGURE 1